Use the compound interest formulas A = P(1 + r/n)^nt and A = Pe^rt to solve.
Find the accumulated value of an investment of $6000 at 7% compounded continuously for 5 years.

Answer by mananth(16949)


$6000 at 7% compounded continuously for 5 years.
P=6000
r= 7%
5 years
continuously
First, convert R as a percent to r as a decimal
r = R/100
r = 7/100
r = 0.07 rate per year,
Then solve the equation for A
A = Pe^rt
A = 6,000.00(2.71828)(0.07)(5)
A = $8,514.41